OpenStudy (anonymous):

Explain the difference between using the sine ratio to solve for a missing angle in a right triangle versus using the cosecant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view

jimthompson5910 (jim_thompson5910):

how are sine and cosecant related?

OpenStudy (anonymous):

isnt it like sin of x is equivalent to cosine of 90-x?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

SOH CAH TOA

OpenStudy (anonymous):

Do you know the ratios?

OpenStudy (anonymous):

?no

jimthompson5910 (jim_thompson5910):

sorry had to step away for a moment

OpenStudy (anonymous):

its cool

OpenStudy (anonymous):

BATMAN SWOOPED IN.....BUT VANISHED

jimthompson5910 (jim_thompson5910):

sine and cosecant are related in this way sin(x) = 1/csc(x)

jimthompson5910 (jim_thompson5910):

oh also csc(x) = 1/sin(x)

jimthompson5910 (jim_thompson5910):

so one is the reciprocal of the other

OpenStudy (anonymous):

so an example would be ..?

jimthompson5910 (jim_thompson5910):

one example is csc(30) = 1/sin(30) here x = 30

OpenStudy (anonymous):

would it be the same if it was sin(30) = 1/csc(30)

jimthompson5910 (jim_thompson5910):

it would be

OpenStudy (anonymous):

thanks again :) ill leave you alone..... for tonight i appreciate it allot :)

jimthompson5910 (jim_thompson5910):

you're welcome